# A Quantum Approach to Synthetic Minority Oversampling Technique (SMOTE)

### v2

@@ Abstract
The paper proposes the Quantum-SMOTE method, a novel solution that uses quantum computing techniques to solve the prevalent problem of class imbalance in machine learning datasets. Quantum-SMOTE, inspired by the Synthetic Minority Oversampling Technique (SMOTE), generates synthetic data points using quantum processes such as swap tests and quantum rotation. The process varies from the conventional SMOTE algorithm’s usage of K-Nearest Neighbors (KNN) and Euclidean distances, enabling synthetic instances to be generated from minority class data points without relying on neighbor proximity. The algorithm asserts greater control over the synthetic data generation process by introducing hyperparameters such as rotation angle, minority percentage, and splitting factor, which allow for customization to specific dataset requirements. The approach is tested on a public dataset of TelecomChurn and evaluated alongside two prominent classification algorithms, Random Forest and Logistic Regression, to determine its impact along with varying proportions of synthetic data.
I Introduction
I-A Unbalanced Classification
@@ I-B Overview of SMOTE
I-C Existing works on SMOTE
@@ I-D Purpose and Scope
I-E Organization
The paper is structured in the following manner. Section II explores the core mathematical principles, including the Basic Concept, several versions of the SMOTE algorithm, and the K-Means Clustering technique. Section III presents an examination of the development of SMOTE utilising quantum techniques, namely the use of the swap test and rotation principles. This is followed by analyzing the outcomes obtained by applying these concepts to actual data. Section IV involves the application of the quantum SMOTE algorithm to a real-world dataset. This process comprises data preparation, clustering, and the production of synthetic data using the SMOTE method. We utilise the SMOTE technique on the telecom data, varying the proportions of the minority class to 30%, 40%, and 50%, respectively. In Section IV-B, we provide a summary of the results and model parameters of the classification Models, which elucidate the effects of Quantum SMOTE.
II Background
II-A Basic Concept of SMOTE
@@ II-E ROC Curve
III Emulating SMOTE Using Quantum
@@
The objective of this strategy is to ensure that the created synthetic data points remain within the statistical bounds of their respective cluster, while also boosting the density of the minority class. In the following sections, we will provide a comprehensive analysis of the algorithm, rotation, and data creation process.
III-A Swap Test
@@
III-B Applying Rotation to data point
@@
III-C Quantum SMOTE Algorithm
@@
IV Case Study and Results
@@
IV-A Improving Telecom Churn Prediction Using SMOTE
@@
IV-A1 Preparing Data For Quantum SMOTE
@@ IV-A2 Clustering
IV-A3 Quantum SMOTE and Synthetic Data
@@
Swap Test: The fundamental operation of the swap test has been previously explained in the preceding sections. We use the swap test in a modified manner to compute the angular distance between the vector representing the minority data point and the vector representing the centroid. The procedure is effectively executed in Ref. [17]. The swap test requires two inputs, denoted as $\phi$ and $\psi$. The state $\phi$ is determined by computing the norms of the inputs, which consist of the centroid and minority data points. On the other hand, the state $\psi$ is obtained by concatenating the normalized components of the inputs. The execution of this preparation is shown in the auxiliary function 3. The circuit that is obtained is rendered in Fig. 6.
@@
Rotation: After performing the swap test, it is necessary to rotate the minority data point by an angle that represents a minute fraction of the total angular distance. The rotation circuit executes the rotation of the normalized minority data point vector. In the preceding section III-B, we have provided a detailed explanation of the different rotations of X, Y, and Z. In this experiment, we applied X rotations to all of the minority data points. To account for numerous interactions or repeated rotations of a single minority data point, we have adjusted the rotation angle by $0.0174$, which corresponds to the conversion from radians to degrees. We are attempting to adjust the angle of the minority data point using angular degrees, even though the angular distance generated by the swap test is in radians. The rotation circuit comprises the state vector of the normalized data point and rotation gates (Fig. 7). By rotating minority data points, synthetic data points that closely resemble the original data points are created, thanks to the use of modest rotation angles. When the synthetic data points are included in the original dataset, it leads to an increase in the total density of the minority class. The scatter distribution of synthetic data points in the population is shown in Fig. 8. The data illustrates the distribution of classes (majority, minority, and synthetic minority) as the proportion of the minority class increases from $30\%$ to $50\%$.
@@
IV-A4 Observation from generated data
Following the generation of synthetic data points by rotation, our next step is to examine the general distribution of important variables throughout the whole population. The objective is to assess if the introduction of artificial data points has caused any significant statistical deviation in the distribution of the variable. The figures 9, 10, 11, and 12 illustrate the distribution of three important variables in the dataset: Tenure, MonthlyCharges, and Total charges. The distribution before the induction of synthetic data points is shown in Fig. 9. The distributions following the induction of synthetic data points, resulting in total minority percentages of $30$, $40$, and $50$, are shown in figures 10, 11, and 12 accordingly. After applying SMOTE, we can confidently state that there is a little distortion to the distribution of variables, but the bins have increased in size. The use of relatively modest angles during rotation prevents any significant deformation to the geometry of the distribution. By comparing the charts depicting the variables after using the SMOTE technique, we see a progressive rise in the values within each category, ranging from $30\%$ to $50\%$. This confirms the successful use of the SMOTE method.
IV-A5 Applying Classification Models
@@
IV-A6 Impact of Quantum SMOTE on Model Statistics
@@
The Random Forest model excels in effectively addressing the Telecom Churn Dataset, particularly when dealing with imbalances in class distribution. The model’s intrinsic advantages, together with its performance improvements using the SMOTE, provide a detailed analysis of its impact in tricky classification scenarios. As we walk through the model’s performance parameters of Confusion matrices (Figs. 15 and 16), Receiver Operating Characteristic Curve (AUC-ROC) (Fig. 17), Precision Recall Curve (AUC) (Fig. 18) we can see gradual improvements with induction of synthetic samples using SMOTE. We discuss the overall improvements in the points below.
@@
Logistic Regression: Performance in the analysis of the Logistic Regression model depicts its ability to handle class imbalance, especially when augmented with the SMOTE. We describe the behavior of Logistic Regression and its outcomes across different scenarios in following sections based on Confusion matrices (Figs. 19 and 20), Receiver Operating Characteristic Curve (AUC-ROC) (Fig. 21), Precision-Recall Curve (AUC) (Fig. 22).
@@
Random Forest
@@
Without Synthetic
1.000 0.784
0.575
0.627 0.811
30% Minority with Synthetic
1.000 0.801
0.634
0.758 0.854
40% Minority with Synthetic
0.996 0.822
0.764
0.888 0.905
50% Minority with Synthetic
0.996 0.846
0.835
0.940 0.929
@@
Without Synthetic
0.797 0.766
0.524
0.604 0.815
30% Minority with Synthetic
0.753 0.759
0.537
0.633 0.812
40% Minority with Synthetic
0.724 0.700
0.608
0.674 0.769
50% Minority with Synthetic
0.732 0.734
0.742
0.779 0.807
TABLE I: Table comparing Accuracy, F1 and AUC score of Random Forest Model for telecom churn dataset without SMOTE, and post SMOTE with minorityas 30%, 40%, and 50%.
Confusion Matrix Comparison
@@
213 181
243
206 387 494 203 774
TABLE II: Table comparing Confusion matrix of Random Forest and Logistic regression without SMOTE, and post SMOTE with minorityas 30%, 40%, and 50%.
IV-A7 Final thoughts on SMOTE Performance

### v3

@@ Abstract
The paper proposes the Quantum-SMOTE method, a novel solution that uses quantum computing techniques to solve the prevalent problem of class imbalance in machine learning datasets. Quantum-SMOTE, inspired by the Synthetic Minority Oversampling Technique (SMOTE), generates synthetic data points using quantum processes such as swap tests and quantum rotation. The process varies from the conventional SMOTE algorithm’s usage of K-Nearest Neighbors (KNN) and Euclidean distances, enabling synthetic instances to be generated from minority class data points without relying on neighbor proximity. The algorithm asserts greater control over the synthetic data generation process by introducing hyperparameters such as rotation angle, minority percentage, and splitting factor, which allow for customization to specific dataset requirements. Due to the use of a compact swap test, the algorithm can accommodate a large number of features. Furthermore, the approach is tested on a public dataset of TelecomChurn and evaluated alongside two prominent classification algorithms, Random Forest and Logistic Regression, to determine its impact along with varying proportions of synthetic data.
keywords:
Quantum-SMOTE, Swaptest, Quantum Rotations
1 Introduction
1.1 Unbalanced Classification
@@ 1.2 Overview of SMOTE
1.3 Existing works on SMOTE
@@ 1.4 Purpose and Scope
1.5 Organization
The paper is structured in the following manner. Section 2 explores the core mathematical principles, including the Basic Concept, several versions of the SMOTE algorithm, and the K-Means Clustering technique. Section 3 presents an examination of the development of SMOTE utilising quantum techniques, namely the use of the swap test and rotation principles. This is followed by analyzing the outcomes obtained by applying these concepts to actual data. Section 4 involves the application of the quantum SMOTE algorithm to a real-world dataset. This process comprises data preparation, clustering, and the production of synthetic data using the SMOTE method. We utilise the SMOTE technique on the telecom data, varying the proportions of the minority class to 30%, 40%, and 50%, respectively. In Section 4.2, we provide a summary of the results and model parameters of the classification Models, which elucidate the effects of Quantum SMOTE.
2 Background
2.1 Basic Concept of SMOTE
@@ 2.5 ROC Curve
3 Emulating SMOTE Using Quantum
@@
The objective of this strategy is to ensure that the created synthetic data points remain within the statistical bounds of their respective cluster while also boosting the density of the minority class. In the following sections, we will provide a comprehensive analysis of the algorithm, rotation, and data creation process.
The figures 1 illustrate fundamental difference in Classical and Quantum SMOTE procedures
Figure 1: Plot illustrating different SMOTE mechanisms. (a) Classical SMOTE, (b) Quantum SMOTE.
3.1 Swap Test
@@
3.1.1 Compact Swaptest
For the purpose of this paper, we have adopted a modified version of the swap test to find the inner product of our two vectors, namely the centroid and an arbitrary minority data point within the cluster. The procedure is already discussed in the articles [17, 18]. Though the article describes the procedure as a dissimilarity measure and uses it to calculate Euclidian distance, we have used it to calculate the inner product of quantum states and thereby the angular distance. The advantage of this procedure is that it requires less number of qubits
$n=\log_{2}(M)+1$
where n is the number of qubits and M is the classical data encoded by amplitude embedding. The procedure is as follows,
We amplitude encode two vectors C (Centroid) and M (Minority) by
$C\longrightarrow|C\rangle$ $=$ $\frac{1}{|C|}\sum_{i}C_{i}\left|q_{i}\right\rangle$ (4)
$M\longrightarrow|M\rangle$ $=$ $\frac{1}{|M|}\sum_{i}M_{i}\left|q_{i}\right\rangle$ (5)
We define the quantum states $|\psi\rangle$ and $|\phi\rangle$ as:
$|\psi\rangle$ $=$ $\frac{|0\rangle\otimes|C\rangle+|1\rangle\otimes|M\rangle}{\sqrt{2}}$
$|\phi\rangle$ $=$ $\frac{|C||0\rangle-|M||1\rangle}{\sqrt{Z}}$
$Z$ $=$ $|C|^{2}+|M|^{2}$ (6)
lets divulge into the details of this circuit
$|0\rangle|\phi\rangle|\psi\rangle$ (7)
$=$ $|+\rangle\left(\frac{(C|0\rangle-M|1\rangle)}{\sqrt{Z}}\right)\left(\frac{|0\rangle|C\rangle+|1\rangle|M\rangle}{\sqrt{2}}\right)$
$=$ $\left(\frac{|0\rangle+|1\rangle}{\sqrt{2}}\right)\left(\frac{C|0\rangle-M|1\rangle}{\sqrt{Z}}\right)\left(\frac{|0\rangle|C\rangle+|1\rangle|M\rangle}{\sqrt{2}}\right)$
$=$ $\frac{1}{2\sqrt{Z}}[|0\rangle(C|0\rangle-M|1\rangle)(|0\rangle|C\rangle+|1\rangle|M\rangle)$
$+|1\rangle(C|0\rangle-M|1\rangle)(|0\rangle|C\rangle+|1\rangle|M\rangle)]$
$=$ $\frac{1}{2\sqrt{Z}}[|0\rangle(C|0\rangle|0\rangle|C\rangle+C|0\rangle|1\rangle|M\rangle-M|1\rangle|0\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle)$
$+|1\rangle(C|0\rangle|0\rangle|C\rangle+C|0\rangle|1\rangle|M\rangle-M|1\rangle|0\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle)]$
Applying controlled swap operation
$=$ $\frac{1}{2\sqrt{Z}}{[|0\rangle(C|0\rangle|0\rangle|C\rangle+C|0\rangle|1\rangle|M\rangle-M|1\rangle|0\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle)}$ (8)
$+$ $|1\rangle(C|0\rangle|0\rangle|C\rangle+C|1\rangle|0\rangle|M\rangle-M|0\rangle|1\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle]$
Applying Hadamard
$=$ $\frac{1}{2\sqrt{Z}}{[|+\rangle(C|0\rangle|0\rangle|C\rangle+C|0\rangle|1\rangle|M\rangle-M|1\rangle|0\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle)}$ (9)
$+$ $|-\rangle(C|0\rangle|0\rangle|C\rangle+C|1\rangle|0\rangle|M\rangle-M|0\rangle|1\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle]$
$=$ $\frac{1}{2\sqrt{2Z}}{[(|0\rangle+|1\rangle)(C|0\rangle|0\rangle|C\rangle+C|0\rangle|1\rangle|M\rangle-M|1\rangle|0\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle)}$
$+$ $(|0\rangle-|1\rangle)(C|0\rangle|0\rangle|C\rangle+C|1\rangle|0\rangle|M\rangle-M|0\rangle|1\rangle|C\rangle-M|1\rangle|1\rangle|M\rangle]$
$=$ $\frac{1}{2\sqrt{2Z}}{[|0\rangle(2C|0\rangle|0\rangle|C\rangle+(C|0\rangle|1\rangle|M\rangle-M|0\rangle|1\rangle|C\rangle)+(C|1\rangle|0\rangle|M\rangle}$
-- $M|1\rangle|0\rangle|C\rangle)-2M|1\rangle|1\rangle|M\rangle)$
$+$ $|1\rangle(C|0\rangle|1\rangle|M\rangle+M|0\rangle|1\rangle|C\rangle-M|1\rangle|0\rangle|C\rangle-C|1\rangle|0\rangle|M\rangle]$
The probability of 0 can be calculated as,
$P_{0}$ $=$ $\frac{1}{8Z}|(2C|0\rangle|0\rangle|C\rangle+(C|0\rangle|1\rangle|M\rangle-M|0\rangle|1\rangle|C\rangle)$
$+$ $(C|1\rangle|0\rangle|M\rangle-M|1\rangle|0\rangle|C\rangle)-2M|1\rangle|1\rangle|M\rangle)|^{2}$
$=$ $\frac{1}{8Z}|(2C|0\rangle|0\rangle|C\rangle+|0\rangle|1\rangle(C|M\rangle-M|C\rangle)$
$+$ $|1\rangle|0\rangle(C|M\rangle-M|C\rangle)-2M|1\rangle|1\rangle|M\rangle)|^{2}$
$=$ $\frac{1}{8Z}|(2C|0\rangle|0\rangle|C\rangle+(|0\rangle|1\rangle+|1\rangle|0\rangle)(C|M\rangle-M|C\rangle)-2M|1\rangle|1\rangle|M\rangle)|^{2}$
$=$ $\frac{1}{8Z}|(2C|0\rangle|0\rangle|C\rangle-2M|1\rangle|1\rangle|M\rangle)+(|0\rangle|1\rangle+|1\rangle|0\rangle)(C|M\rangle-M|C\rangle)|^{2}$
$=$ $\frac{1}{8Z}(|2C|0\rangle|0\rangle|C\rangle-2M|1\rangle|1\rangle|M\rangle|^{2}+||0\rangle|1\rangle+|1\rangle|0\rangle|^{2}|C|M\rangle-M|C\rangle|^{2})$
$=$ $\frac{1}{8Z}(4C^{2}+4M^{2}+||0\rangle|1\rangle+|1\rangle|0\rangle|^{2}|C|M\rangle-M|C\rangle|^{2})$
$=$ $\frac{1}{8Z}(4Z+2(C^{2}+M^{2}-2CM\braket{M}{C}))$
$=$ $\frac{1}{8Z}(4Z+2(Z-2CM\braket{M}{C}))$
$=$ $\frac{1}{4Z}(2Z+(Z-2CM\braket{M}{C}))$
$=$ $\frac{1}{4Z}(3Z-2CM\braket{M}{C})$
$\implies\braket{M}{C}$ $=$ $\frac{(3-4P_{0})Z}{2CM}$ (10)
The above equation 3.1.1 states that after measurement, from the probability of 0, we obtain the inner product between the centroid and minority. In a slightly different perspective, let us calculate inner product of $\psi$ and $\phi$,
$\langle\phi\mid\psi\rangle=\left(\frac{\langle C|\otimes\langle 0|-\langle M|\otimes\langle 1|}{\sqrt{Z}}\right)\left(\frac{|0\rangle\otimes|C\rangle+|1\rangle\otimes|M\rangle}{\sqrt{2}}\right)$ (11)
Expanding the inner product:
$\langle\phi\mid\psi\rangle$ $=$ $\frac{1}{\sqrt{Z}}\frac{1}{\sqrt{2}}(\langle C|\otimes\langle 0|(|0\rangle\otimes|C\rangle)+\langle C|\otimes\langle 0|(|1\rangle\otimes|M\rangle)$ (12)
-- $\langle M|\otimes\langle 1|(|0\rangle\otimes|C\rangle)-\langle M|\otimes\langle 1|(|1\rangle\otimes|M\rangle))$
Simplifying each term:
$\langle C|\otimes\left\langle\left.0|(|0\rangle\otimes|C\rangle)=\langle C\mid C\rangle\otimes\langle 0\mid 0\rangle=|C\right|^{2}\right.$
$\langle C|\otimes\langle 0|(|1\rangle\otimes|M\rangle)=0$
$\langle M|\otimes\langle 1|(|0\rangle\otimes|C\rangle)=0$
$\langle M|\otimes\left\langle\left.1|(|1\rangle\otimes|M\rangle)=\langle M\mid M\rangle\otimes\langle 1\mid 1\rangle=|M\right|^{2}\right.$ (13)
So, the inner product simplifies to:
$\langle\phi\mid\psi\rangle$ $=$ $\frac{1}{\sqrt{Z}}\frac{1}{\sqrt{2}}\left(|C|^{2}-|M|^{2}\right)$
$\langle\phi\mid\psi\rangle$ $=$ $\frac{|C|^{2}-|M|^{2}}{\sqrt{2Z}}$ (14)
Calculating $|\langle\phi\mid\psi\rangle|^{2}$ :
$|\langle\phi\mid\psi\rangle|^{2}=\left(\frac{|C|^{2}-|M|^{2}}{\sqrt{2Z}}\right)^{2}=\frac{\left(|C|^{2}-|M|^{2}\right)^{2}}{2Z}$ (15)
${2Z}|\langle\phi\mid\psi\rangle|^{2}=2Z\left(\frac{\left(|C|^{2}-|M|^{2}\right)^{2}}{2Z}\right)$ (16)
simplifying:
${2Z}|\langle\phi\mid\psi\rangle|^{2}=\left(|C|^{2}-|M|^{2}\right)^{2}$ (17)
Assuming
${2Z}|\langle\phi\mid\psi\rangle|^{2}$ $=$ $D^{2}$
$\implies D^{2}$ $=$ $2Z|\langle\phi\mid\psi\rangle|^{2}$ (18)
The term $D$ refers to the euclidean distance [18], and the inner product of $\langle\phi\mid\psi\rangle$ represents the swaptest probability.
[…]
Figure 3: Compact Swap test circuit.
@@ 3.1.1 Compact Swaptest
3.2 Applying Rotation to data point
@@
3.3 Quantum SMOTE Algorithm
@@
4 Case Study and Results
@@
4.1 Improving Telecom Churn Prediction Using SMOTE
@@
4.1.1 Preparing Data For Quantum SMOTE
@@ 4.1.2 Clustering
4.1.3 Quantum SMOTE and Synthetic Data
@@
Swap Test: The fundamental operation of the swap test has been previously explained in the preceding sections. We use the swap test in a modified manner 3.1.1 to compute the angular distance between the vector representing the minority data point and the vector representing the centroid. The procedure is effectively executed in Ref. [17, 18]. The swap test requires two inputs, denoted as $\phi$ and $\psi$. The state $\phi$ is determined by computing the norms of the inputs, which consist of the centroid and minority data points. On the other hand, the state $\psi$ is obtained by concatenating the normalized components of the inputs. The execution of this preparation is shown in the auxiliary function 3. The circuit that is obtained is rendered in Fig. 3.
@@
Rotation: After performing the swap test, it is necessary to rotate the minority data point by an angle that represents a minute fraction of the total angular distance. The rotation circuit executes the rotation of the normalized minority data point vector. In the preceding section 3.2, we have provided a detailed explanation of the different rotations of X, Y, and Z. In this experiment, we applied X rotations to all of the minority data points. To account for numerous interactions or repeated rotations of a single minority data point, we have adjusted the rotation angle by $0.0174$, which corresponds to the conversion from radians to degrees. We are attempting to adjust the angle of the minority data point using angular degrees, even though the angular distance generated by the swap test is in radians. The rotation circuit comprises the state vector of the normalized data point and rotation gates (Fig. 8). By rotating minority data points, synthetic data points that closely resemble the original data points are created, thanks to the use of modest rotation angles. When the synthetic data points are included in the original dataset, it leads to an increase in the total density of the minority class. The scatter distribution of synthetic data points in the population is shown in Fig. 9. The data illustrates the distribution of classes (majority, minority, and synthetic minority) as the proportion of the minority class increases from $30\%$ to $50\%$.
@@
4.1.4 Observation from generated data
Following the generation of synthetic data points by rotation, our next step is to examine the general distribution of important variables throughout the whole population. The objective is to assess if the introduction of artificial data points has caused any significant statistical deviation in the distribution of the variable. The figures 10, 11, 12, and 13 illustrate the distribution of three important variables in the dataset: Tenure, MonthlyCharges, and Total charges. The distribution before the induction of synthetic data points is shown in Fig. 10. The distributions following the induction of synthetic data points, resulting in total minority percentages of $30$, $40$, and $50$, are shown in figures 11, 12, and 13 accordingly. After applying SMOTE, we can confidently state that there is a little distortion to the distribution of variables, but the bins have increased in size. The use of relatively modest angles during rotation prevents any significant deformation to the geometry of the distribution. By comparing the charts depicting the variables after using the SMOTE technique, we see a progressive rise in the values within each category, ranging from $30\%$ to $50\%$. This confirms the successful use of the SMOTE method.
4.1.5 Applying Classification Models
@@
4.1.6 Impact of Quantum SMOTE on Model Statistics
@@
The Random Forest model excels in effectively addressing the Telecom Churn Dataset, particularly when dealing with imbalances in class distribution. The model’s intrinsic advantages, together with its performance improvements using the SMOTE, provide a detailed analysis of its impact in tricky classification scenarios. As we walk through the model’s performance parameters of Confusion matrices (Figs. 16 and 17), Receiver Operating Characteristic Curve (AUC-ROC) (Fig. 18), Precision Recall Curve (AUC) (Fig. 19) we can see gradual improvements with induction of synthetic samples using SMOTE. We discuss the overall improvements in the points below.
@@
Logistic Regression: Performance in the analysis of the Logistic Regression model depicts its ability to handle class imbalance, especially when augmented with the SMOTE. We describe the behavior of Logistic Regression and its outcomes across different scenarios in following sections based on Confusion matrices (Figs. 20 and 21), Receiver Operating Characteristic Curve (AUC-ROC) (Fig. 22), Precision-Recall Curve (AUC) (Fig. 23).
@@
{tblr}
width = colspec = Q[377]Q[117]Q[117]Q[138]Q[88]Q[88], row1 = Parsley,c,fg=white, row2 = c, row8 = RoyalPurple,c,fg=white, column3 = c, column6 = c, cell11 = c=60.924, cell22 = c=20.234, cell25 = c=20.176, cell32 = c, cell34 = c, cell35 = c, cell42 = c, cell43 = TitanWhite, cell44 = MintTulip,c, cell45 = WePeep,c, cell46 = Concrete, cell52 = c, cell53 = Skeptic, cell54 = BlizzardBlue,c, cell55 = LightOrchid,c, cell56 = Texas, cell62 = c, cell63 = MossGreen, cell64 = TurquoiseBlue,c, cell65 = Orchid,c, cell66 = GoldenFizz, cell72 = c, cell73 = Fern, cell74 = PictonBlue,c, cell75 = Orchid1,c, cell76 = yellow, cell81 = c=60.924, cell92 = c, cell93 = Fern, cell94 = MintTulip,c, cell95 = WePeep,c, cell96 = yellow, cell102 = c, cell103 = DeYork, cell104 = BlizzardBlue1,c, cell105 = WePeep1,c, cell106 = Yellow, cell112 = c, cell113 = TitanWhite, cell114 = Cornflower,c, cell115 = LightOrchid1,c, cell116 = Concrete, cell122 = c, cell123 = FringyFlower, cell124 = PictonBlue,c, cell125 = Orchid1,c, cell126 = GoldenFizz1, vlines, hline1-2,4-13 = -, hline3 = 1-3,5-6, Random Forest &
@@
Without Synthetic 1.000 0.784 0.575 0.627 0.811
30% Minority with Synthetic 1.000 0.801 0.634 0.758 0.854
40% Minority with Synthetic 0.996 0.822 0.764 0.888 0.905
50% Minority with Synthetic 0.996 0.846 0.835 0.940 0.929
@@
Without Synthetic 0.797 0.766 0.524 0.604 0.815
30% Minority with Synthetic 0.753 0.759 0.537 0.633 0.812
40% Minority with Synthetic 0.724 0.700 0.608 0.674 0.769
50% Minority with Synthetic 0.732 0.734 0.742 0.779 0.807
Table 1: Table comparing Accuracy, F1 and AUC score of Random Forest Model for telecom churn dataset without SMOTE, and post SMOTE with minorityas 30%, 40%, and 50%.
{tblr}
width = colspec = Q[248]Q[90]Q[90]Q[79]Q[79]Q[86]Q[69]Q[79]Q[79], cells = c, row1 = MidnightBlue,fg=white, row3 = PastelGreen, row5 = PastelGreen1, row7 = WineBerry, row8 = GrannySmithApple, row10 = GrannySmithApple, cell11 = c=90.898, cell21 = r=5Cornflower,fg=FreshEggplant, cell22 = c=20.18, cell24 = c=20.158, cell26 = c=20.155, cell28 = c=20.158, cell32 = Malibu, cell33 = Malibu, cell36 = Tacao, cell37 = Tacao, cell38 = Turquoise, cell39 = Turquoise, cell52 = Malibu, cell53 = Malibu, cell56 = Tacao, cell57 = Tacao, cell58 = Turquoise, cell59 = Turquoise, cell71 = r=5WePeep,fg=blue, cell72 = c=80.651, cell82 = LightOrchid, cell83 = LightOrchid, cell86 = Manhattan, cell87 = Manhattan, cell88 = BrightTurquoise, cell89 = BrightTurquoise, cell102 = LightOrchid, cell103 = LightOrchid, cell106 = Manhattan, cell107 = Manhattan, cell108 = BrightTurquoise, cell109 = BrightTurquoise, vlines, hline1-2,7,12 = -, hline3-6,8-11 = 2-9, Confusion Matrix Comparison &
@@
213 181 243 206 387 494 203 774
Table 2: Table comparing Confusion matrix of Random Forest and Logistic regression without SMOTE, and post SMOTE with minorityas 30%, 40%, and 50%.
4.1.7 Final thoughts on SMOTE Performance